OpenStudy (anonymous):

evaluate each expression if a=4 b=-10 and c=-4 3a+4b+6c

OpenStudy (anonymous):

3a +4b+6c = 3(4) +4(-10) +6(-4) = 12-40-24 =-52

OpenStudy (anonymous):

thanks